### v5

# Global Geolocated Realtime Data of Interfleet Urban Transit Bus Idling

Abstract
Urban transit bus idling is a contributor to ecological stress, economic inefficiency, and medically hazardous health outcomes due to emissions. The global accumulation of this frequent pattern of undesirable driving behavior is enormous. In order to measure its scale, we propose GRD-TRT-BUF-4I (“Ground Truth Buffer for Idling”) an extensible, realtime detection system that records the geolocation and idling duration of urban transit bus fleets internationally. Using live vehicle locations from General Transit Feed Specification (GTFS) Realtime, the system detects approximately 200,000 idling events per day from over 50 cities across North America, Europe, Oceania, and Asia. This realtime data was created dynamically to serve operational decision-making and fleet management to reduce the frequency and duration of idling events as they occur, as well as to capture its accumulative effects. Civil and Transportation Engineers, Urban Planners, Epidemiologists, Policymakers, and other stakeholders might find this useful for emissions modeling, traffic management, route planning, and other urban sustainability efforts at a variety of geographic and temporal scales.
1 Introduction
Urban transit buses provide essential transportation services in cities around the world. In Brazil, they are the most common mode of transportation [1]. In India, they are the third most preferred mode of transportation, where the size of their national fleet increased by nearly 30% between 2009 and 2019, from nearly 118,00 vehicles to more than 152,000 [2, 3]. China also increased the size of its national fleet over roughly the same period, totaling nearly 600,000 vehicles in 2021 [4]. In the United States, nearly half of all passenger trips on public transportation were facilitated by urban transit buses, amounting to approximately 6.2 billion trips in 2022, most of which were powered by diesel fuel [5, 6].
Data from the American Public Transportation Association (APTA) reported that in 2020, more than 96% of all commuter buses in the United States were diesel-powered [6]. All other fuel types such as compressed natural gas (CNG), liquefied natural gas (LNG), biodiesel, gasoline, propane, hydrogen, hybrid, and battery-electric accounted for less than 4% of all fielded vehicles when combined [6]. Even throughout Europe, where aggressive measures have been taken to transition away from diesel-powered vehicles, the most common type of newly fielded urban transit bus in 2021 was indeed diesel-powered [7]. Although many implications emerge in this regard, engine idling is a concern worldwide.
Idling is when a vehicle’s engine is in operation, but is geographically stationary [8]. While individual idling events may seem insignificant, the total accumulation of these ephemeral, but repeated events is material. For instance, in 2022 the National Renewable Energy Laboratory (NREL) found that urban transit buses in the United States idled for roughly 40% of their typical 9-hour operational period based on data from 16 vehicles over 19,440 hours (2.22 years) of cumulative driving time [9]. An earlier study from 2010 reported similar findings throughout the United Kingdom, where idling rates were between 30% and 44% [10]. A smaller scale, but more detailed example in South Korea found similar idling rates near 43% [11].
Ecological stress, economic inefficiency, and medically hazardous health outcomes - among others - all simultaneously emerge from this frequent and repeated pattern of undesirable driving behavior [12, 13, 14]. A single idling urban transit bus can consume 0.5 gal (1.9 L) of diesel fuel per hour, emitting approximately 11.1 lbs (5 kg) of $\text{CO}_{2}$ equivalent GHG, 2.7 g of VOC, 2.7 g of THC, 37.4 g of CO, 61.1 g of $\text{NO}_{x}$, 1.2 g of $\text{PM}_{10}$, and 1.1 g of $\text{PM}_{2.5}$ [15, 16]. These pollutants pose serious health risks including asthma, cardiovascular disease, liver and kidney damage [17, 18, 19]. Furthermore, noise from urban transit bus idling is a reported burden loud enough to potentially cause permanent hearing loss at SEL and $L_{max}$ A-weighted decibel levels of 111 dBA and 75 dBA, respectively [13, 14].
While the negative impacts of urban transit bus idling are numerous and evident, the magnitude of the problem has begun to emerge as a category of its own concern [20, 21]. Existing studies tend to focus on historical information from selected localities that are limited to a municipal, regional, or national fleet [12, 22, 23]. Few if any comprehensively measure the interfleet phenomena in realtime on an international scale. Well-established developments like the General Transit Feed Specification (GTFS) Realtime have enabled transit agencies to share realtime data on service disruptions, vehicle locations, and arrival times using a standard schema since 2011 [24, 25]. Growing international adoption of GTFS Realtime now affords the possibility to integrate these data sources using a common ingestion method [21], as exhibited here.
This geolocated realtime data is the first record of urban transit bus idling measured on a global scale. It immediately describes the worldwide phenomena when and where it occurs, giving transit agencies around the world the collective ability to actively intercept idling events, rather than merely measure the problem, only to analyze its negative impact after the fact. This shift would mark a significant advancement in the approach to reducing urban transit bus fleet idling worldwide by moving from reactive policies based on historical trends to preemptive measures based on proactive reduction strategies. The extensibility of this effort also allows easy integration of new GTFS Realtime sources as they become available. Those that fit within this scope are described in the following section Data Sources.
2 Methods
2.1 Data Sources
Live vehicle locations from on-network urban transit bus fleets from over 50 cities were collected using GTFS Realtime. Although there are hundreds of GTFS Realtime sources, those that were selected, satisfied a combination of immediate availability, geographic variety, sufficient fleet size, daily operational length, and active support of each endpoint. Note that transit agencies may not elect to adopt GTFS Realtime for reasons regarding development cost, technical expertise and maintenance, usage of existing systems, lack of political support or priority [26, 27].
To ensure authenticity and reproducibility, data from GTFS Realtime was collected from publicly accessible Representational State Transfer (REST) Application Programming Interface (API) endpoints provided by the transit agencies directly, rather than relying on private third-party aggregation. All data sources were consumed as Protocol Buffers, sometimes referred to as “protobufs” [28, 29]. An example of the deserialized protobuf schema is exhibited in Figure 2 and Figure 3.
Each data source was categorized according to its geographic region, where each region contained between 1 to 6 sources depending on the availability and stability of the REST API. Table 2 through Table 12 exhibit each data source and its corresponding region. Some regions were nested within national categories belonging to the same continent. In other cases, few enough data sources were available that the continent directly represented the region, such as Oceania and Asia.
Refer to captionNorth AmericaEuropeOceaniaAsia
Figure 1: Geographic Distribution of GTFS Realtime Sources by Region [30]
Continent Country Region
North America United States US East
US West
US South
US Central
US Mountain
Canada Canada East
Canada West
Europe Greater EU West
EU Central
Oceania Greater Oceania
Asia Greater Asia
Table 1: GTFS Realtime Sources by Region
Greater indicates the corresponding region contains more than one country.
Note that in Table 1, the Greater designation in the Country column indicates the corresponding region contains more than one country. Similarly, the italicized names in the City column in Table 2 through Table 12 include neighboring cities from the broader metropolitan area. For example, Minneapolis includes neighboring St. Paul, San Francisco includes neighboring Oakland, etc. Other cities may also include locations beyond their proper municipal geographic boundary. However, those that are stylized were specified as such because of their immediate availability.
Region City Agency Ref.
US East New York Metropolitan Transportation Authority (MTA) [31]
Philadelphia Southeastern Pennsylvania Transport. Authority (SEPTA) [32]
Wash. D.C. Washington Metropolitan Area Transit Authority (WMATA) [33]
Boston Massachusetts Bay Transportation Authority (MBTA) [34]
Pittsburgh Pittsburgh Regional Transit (PRT) [35]
Table 2: US East GTFS Realtime Sources
Italicized city names include the broader metropolitan area.
Region City Agency Ref.
US West Los Angeles Los Angeles County Metro. Transport. Authority (Metro) [36]
San Fran. Metropolitan Transportation Commission (MTC) [37]
San Diego San Diego Metropolitan Transit System (MTS) [38]
Seattle King County Metro (KCM) [39]
Sacramento Sacramento Regional Transit District (SacRT) [40]
Portland Tri-County Metro. Transport. District of Oregon (TriMet) [41]
Table 3: US West GTFS Realtime Sources
Italicized city names include the broader metropolitan area.
Region City Agency Ref.
US South Atlanta Metropolitan Atlanta Rapid Transit Authority (MARTA) [42]
Miami Miami-Dade Transit (MDT) [43]
Tampa Hillsborough Area Regional Transit (HART) [44]
Louisville Transit Authority of River City (TARC) [45]
Nashville Nashville Metropolitan Transit Authority (Nashville MTA) [46]
Table 4: US South GTFS Realtime Sources
Region City Agency Ref.
US Central Minneapolis Metro Transit (Minnesota) [47]
St. Louis Metro Transit (St. Louis) [48]
Madison Metro Transit (Madison) [49]
Columbus Central Ohio Transit Authority (COTA) [50]
Des Moines Des Moines Area Regional Transit Authority (DART) [51]
Table 5: US Central GTFS Realtime Sources
Italicized city names include the broader metropolitan area.
Region City Agency Ref.
US Mountain Denver Regional Transportation District (RTD) [52]
Phoenix Valley Metro Reg. Pub. Transport. Auth. (Valley Metro) [53]
San Antonio VIA Metropolitan Transit Authority (VIA Metro) [54]
Austin Capital Metropolitan Transport. Authority (CapMetro) [55]
Billings Billings Metropolitan Transit (MET) [56]
Table 6: US Mountain GTFS Realtime Sources
Region City Agency Ref.
Canada East Montréal Société de transport de Montréal (STM) [57]
York York Region Transit (YRT) [58]
Hamilton Hamilton Street Railway (HSR) [59]
Halifax Halifax Transit [60]
Thunder Bay Thunder Bay Transit [61]
Table 7: Canada East GTFS Realtime Sources
Region City Agency Ref.
Canada West Vancouver TransLink (British Columbia) [62]
Calgary Calgary Transit [63]
Edmonton Edmonton Transit System (ETS) [64]
Saskatoon Saskatoon Transit [65]
Table 8: Canada West GTFS Realtime Sources
Region City Agency Ref.
EU West Amsterdam Openbaar Vervoer (OV) [66]
Stockholm Storstockholms Lokaltrafik (SL) [67]
Helsinki Helsinki Regional Transport Authority (HSL) [68]
Dublin National Transport Authority [69]
Rome Azienda Tramvie e Autobus del Comune di Roma (ATAC) [70]
Table 9: EU West GTFS Realtime Sources
Italicized city names include the broader country.
Region City Agency Ref.
EU Central Warsaw Zarząd Transportu Miejskiego w Warszawie (ZTM) [71]
Krakow Zarząd Transportu Publicznego w Krakowie (ZTP) [72]
Gdansk Zarządu Transportu Miejskiego w Gdańsku (ZTM) [73]
Prague Pražská Integrovaná Doprava (PID) [74]
Table 10: EU Central GTFS Realtime Sources
Region City Agency Ref.
Oceania Sydney Transport for New South Wales (Transport for NSW) [75]
Brisbane Translink (Queensland) [76]
Adelaide Adelaide Metro [77]
Auckland Auckland Transport (AT) [78]
Christchurch Environment Canterbury (Metro) [79]
Table 11: Oceania GTFS Realtime Sources
Region City Agency Ref.
Asia Delhi Delhi Transport Corporation (DTC) [80]
Table 12: Asia GTFS Realtime Sources
gtfs_realtime_version: "2.0" incrementality: FULL_DATASET timestamp: 1706573497
Figure 2: Example of GTFS Realtime Protocol Buffer Header Object
[ id: "y0811" vehicle { trip { trip_id: "60487628" route_id: "216" direction_id: 1 start_time: "19:16:00" start_date: "20240129" schedule_relationship: SCHEDULED } vehicle { id: "y0811" label: "0811" } position { latitude: 42.2721062 longitude: -70.9509277 bearing: 0 } current_stop_sequence: 1 stop_id: "3265" current_status: STOPPED_AT timestamp: 1706573492 occupancy_status: MANY_SEATS_AVAILABLE occupancy_percentage: 0 } ]
Figure 3: Example of GTFS Realtime Protocol Buffer Message Object
2.2 Buffering Procedure
Each geographic region (US East , US West , etc.) previously mentioned in Data Sources, represents an asynchronous feed, $\text{Feed}_{i}$, exhibited in Figure 4. It is constructed from its corresponding GTFS Realtime sources. As $\text{Feed}_{i}$ enters the buffer, it is stored in memory as $d_{i}$ according to indices $d_{0},d_{1},d_{2},$$d_{h+1}$ at time $t,t+1,t+2,$,$t+h+1,$ respectively.
After the buffer is populated to length $T$, the buffering procedure is fully initialized. As new $\text{Feed}_{i}$ enters the buffer at time $t$ once initialized, $\text{Feed}_{i+h+2}$ is omitted beyond buffer length $T$. This step is repeated theoretically to $\infty$ or practically until the procedure is stopped.
After initializing the buffer, as well as when $\text{Feed}_{i}$ enters it, the index $d_{0}$ is treated as set $A$, index $d_{h}$ treated as set $B$, and index $d_{h+1}$ treated as set $C$. In other words, sets $A,B,C$ are equivalent to buffer indices $d_{0},d_{h},d_{h+1}$, respectively. The subsetting procedure is fully initialized with sets $A,B,C$.
$n(A\cap B)$ is then computed, such that $A^{n}$ and $B^{n}$ are of length $n$. $\mathcal{A}$ and $\mathcal{B}$ are also of length $n$ and are further subsets of $A$ and $B$, only containing elements $x$ and $y$, respectively, subset from attributes $j$. Similarly, elements $z$ in $\mathcal{C}$ are subset from the same attributes $j$. $\mathcal{H}_{i}$ is then computed from $\mathcal{A}\cap\mathcal{B}$. Finally, $\mathcal{Y}=\mathcal{H}\cap\mathcal{C}$. The following section Subsetting Procedure provides more detail regarding the time-series set approach.
[…]
Figure 4: GTFS Realtime Buffering & Subsetting Procedure
2.3 Subsetting Procedure
Recall from Buffering Procedure that $A$ is a set, and assume $a$ is a positive integer denoting the tuples whose elements $x_{(i,j)}$ at time $t$ belong to $A$, denoted as $A^{a}$, such that:
$A^{a}=\{x_{(i,j)\,t}:\text{for all}\,x\in A\}$ (1)
when expanded by $i$:
$A^{a}=\Biggl{\{}\begin{bmatrix}x_{(1,j)\,t}\\ x_{(2,j)\,t}\\ \vdots\\ x_{(a,j)\,t}\end{bmatrix}:\text{for all}\,x\in A\Biggl{\}}$ (2)
and further by $j$:
$A^{a}=\bigl{\{}\begin{bmatrix}x_{(i,j=1,2,\ldots,m)\,t}\\ \end{bmatrix}:\text{for all}\,x\in A\bigl{\}}$ (3)
where:
[…]
and $m$ is a positive integer denoting the attribute size.
Similarly, $B$ is a set, and $b$ is a positive integer denoting the tuples whose elements $y_{(i,j)}$ at time $t+h$ belong to $B$, denoted as $B^{b}$, such that:
$B^{b}=\{y_{(i,j)\,t+h}:\text{for all}\,y\in B\}$ (5)
where:
[…]
The intersection of $A^{a}$ and $B^{b}$ is then computed as:
$n(A^{a}\cap B^{b})=A^{a}\cap B^{b}$ (7)
where subsets $A^{n}$ and $B^{n}$ now have have equal cardinality and $n$ denotes the equal index length, such that $|A^{n}|=|B^{n}|$, where:
$\begin{aligned} A^{n}&=\{x_{(i,j)\,t}:x\in A\}\\ \end{aligned}\qquad\begin{aligned} B^{n}&=\{y_{(i,j)\,t+h}:y\in B\}\end{aligned}$ (8)
when expanded by the $i^{th}$ index:
$\begin{aligned} A^{n}=\Biggl{\{}\begin{bmatrix}x_{(1,j)\,t}\\ x_{(2,j)\,t}\\ \vdots\\ x_{(n,j)\,t}\end{bmatrix}:x\in A\Biggl{\}}\end{aligned}\qquad\begin{aligned} B^{n}=\Biggl{\{}\begin{bmatrix}y_{(1,j)\,t+h}\\ y_{(2,j)\,t+h}\\ \vdots\\ y_{(n,j)\,t+h}\end{bmatrix}:y\in B\Biggl{\}}\end{aligned}$ (9)
Further denoting subsets $A^{n}\subseteq A^{b}$ and $B^{n}\subseteq B^{b}$, we have:
$\begin{aligned} \mathcal{A}&=\bigl{\{}\begin{bmatrix}x_{(i,j=1,\ldots,5)\,t}\\ \end{bmatrix}:x\in A^{n}\bigl{\}}\end{aligned}\qquad\begin{aligned} \mathcal{B}&=\bigl{\{}\begin{bmatrix}y_{(i,j=1,\ldots,5)\,t+h}\\ \end{bmatrix}:y\in B^{n}\bigl{\}}\end{aligned}$ (10)
so that attributes $j=1,\ldots,5$ in subsets $A^{n}$ and $B^{n}$ are referred to directly as $\mathcal{A}$ and $\mathcal{B}$ from $t$ to $t+h$, respectively, and where $j=1,\ldots,5$ is:
$\begin{split}1&=\texttt{vehicle\_id}\\ 2&=\texttt{route\_id}\\ 3&=\texttt{trip\_id}\\ 4&=\texttt{latitude}\\ 5&=\texttt{longitude}\end{split}$ (11)
Next, the subset $\mathcal{H}$ is computed as the intersection of $\mathcal{A}$ and $\mathcal{B}$, where $\mathcal{H}\subseteq\mathcal{A}$ and $\mathcal{H}\subseteq\mathcal{B}$, whose elements $w_{(i,j)}$ at time $t+h$ belong to $\mathcal{H}$, such that:
$\mathcal{H}=\bigl{\{}\begin{bmatrix}w_{(i;j)\,t+h}\\ \end{bmatrix}:w\in\mathcal{A}\cap\mathcal{B}\}$ (12)
when expanded by both $i^{th}$ index and $j^{th}$ attributes:
$\mathcal{H}^{p}=\Biggl{\{}\begin{bmatrix}w_{(1;\,j=1,\ldots,5)\,t+h}\\ w_{(2;\,j=1,\ldots,5)\,t+h}\\ \vdots\\ w_{(p;\,j=1,\ldots,5)\,t+h}\end{bmatrix}:w\in\mathcal{A}\cap\mathcal{B}\Biggl{\}}$ (13)
It is important to highlight that the subset $\mathcal{H}$ is treated as a special case whose elements $w_{(i,j)}$ can remain in the subset throughout any sequence of time-steps and is appended as $\mathcal{H}_{i}$ to length $\mathcal{H}^{p}$. This differs from $\mathcal{A}$ and $\mathcal{B}$ whose elements can change at every time-step.
Similar to original sets $A$ and $B$, let $C$ be a set, and $c$ be a positive integer denoting the tuples whose elements $z_{(i,j)}$ at time $t+h+1$ belong to set $C$, denoted as $C^{c}$. This differs from the convention that $c$ is the complement, such that:
$C^{c}=\{z_{(i,j)\,t+h+1}:\text{for all}\,z\in C\}$ (14)
when expanded by the $i^{th}$ index:
$C^{c}=\Biggl{\{}\begin{bmatrix}z_{(1,j)\,t+h+1}\\ z_{(2,j)\,t+h+1}\\ \vdots\\ z_{(c,j)\,t+h+1}\end{bmatrix}:\text{for all}\,z\in C\Biggl{\}}$ (15)
To match the attributes in set $C$ to those found in subsets $\mathcal{A}$ and $\mathcal{B}$, let the following be true:
$\mathcal{C}$ $=\bigl{\{}\begin{bmatrix}{}_{(i,j=1,\ldots,5)\,t+h+1}\\ \end{bmatrix}:z\in C\bigl{\}}$ (16)
so that attributes $j=1,\ldots,5$ in set $C$ are referred to directly as subset $\mathcal{C}$ at time $t+h+1$. Next, the intersection of $\mathcal{C}$ and $\mathcal{H}$ is computed, where:
$\mathcal{Y}=\mathcal{C}\cap\mathcal{H}$ (17)
$\mathcal{Y}$ is now a further subset of $\mathcal{C}$ and $\mathcal{H}$, $\mathcal{Y}\subseteq\mathcal{C}$ and $\mathcal{Y}\subseteq\mathcal{H}$, whose elements $v_{(i,j)}$ at time $T$ belong to $\mathcal{Y}$, such that:
$\mathcal{Y}=\bigl{\{}\begin{bmatrix}v_{(i;j)\,T}\\ \end{bmatrix}:v\in\mathcal{C}\cap\mathcal{H}\}$ (18)
when expanded by both $i^{th}$ index and $j^{th}$ attributes:
$\mathcal{Y}^{k}=\Biggl{\{}\begin{bmatrix}v_{(1,j=1,\ldots,5)\,T}\\ v_{(2,j=1,\ldots,5)\,T}\\ \vdots\\ v_{(k,j=1,\ldots,5)\,T}\end{bmatrix}:v\in\mathcal{C}\cap\mathcal{H}\Biggl{\}}$ (19)
Note that the subset $\mathcal{Y}$ is also treated as a special case whose elements $v_{(i,j)}$ can remain in the subset throughout any sequence of time-steps. However, it is not appended, as it is the final subset of length $\mathcal{Y}^{k}$. As the final output, $\mathcal{Y}$ resolves this series of expressions describing the subsetting procedure. The following section Computational Procedure describes its parameters and how it’s computed.
2.4 Computational Procedure
Combining Buffering Procedure and Subsetting Procedure, Algorithm 1 GRD-TRT-BUF-4I (“Ground Truth Buffer for Idling") casts them into step-by-step instructions for computing realtime idling events. Details of the algorithm are provided in Appendix. There are three parameters:
  1. 1.
    $r$: the rate at which GTFS Realtime server requests are made (seconds). It is a positive integer. $r$ has a default value of 30 seconds and should not exceed the rate limit of any GTFS Realtime server. It is recommended that $r$ not take values less than the frequency of vehicle location updates, which are typically updated every 30 seconds.
  2. 2.
    $h$: the number of time-steps $t$ after an idling event is measured (interval). It is a positive integer. $h$ has a default value of 1. In the case that the default value of $r$ remains unchanged, any number of seconds beyond 60 seconds is recorded as an idling event. Alternatively, specifying $h$ as 2 would consider an idling event any number of seconds beyond 90 seconds, all else equal.
  3. 3.
    $m$: a constant used to bound the length of appended subset $\mathcal{H}$ (iterations). It is a positive integer. $m$ has an arbitrary default value of 10. It is the maximum allowable iterations that elements $z\in\mathcal{H}$ are stored before being omitted when also not found in set $\mathcal{C}$. In other words, the elements $z$ are removed from the appended set $\mathcal{H}$ after $m$ number of iterations they are not within $\mathcal{H}\cap C$.
After $r$, $h$, and $m$ are specified, an infinite loop is conducted outside the process in which Buffering Procedure and Subsetting Procedure are both initialized and computed. Additional instructions in GRD-TRT-BUF-4I from steps 10 to 22 correspond to the subset $\mathcal{H}$. Again, details of the algorithm are provided in Appendix.
Finally, $\mathcal{Y}$ is computed as the primary output and is immediately returned and stored on disk. The following section System Architecture & Design broadly outlines a collection of components and their relationships in this regard.
2.5 System Architecture & Design
2.5.1 System Architecture
The system contains the basic functions of Algorithm 1 GRD-TRT-BUF-4I. It is modeled through the Integration Definition for Process Modeling (IDEF0), exhibited in Figure 5. Each component takes an input (left pointer), yields an output (right pointer), and requires a mechanism (bottom pointer). Those that process intermediate results also take control parameters (top pointer).
First, the Serve component inputs “Fleet” - the telemetry data collected from the live location of vehicles in the network of urban transit bus fleets. It outputs $\text{Feed}_{i}$”, the GTFS Realtime feeds, and requires the open internet or “Web” as a mechanism. Note that this process is modeled for architectural context, where no explicit controls are required.
The Extract component inputs the GTFS Realtime sources, $\text{Feed}_{i}$ and outputs the extracted data $d_{i}$, requiring an API key(s) or “Key” as a mechanism to authorize REST API endpoint(s) access where required. The request rate is controlled by $r$, as previously described in Computational Procedure. The purpose of this component is to asynchronously extract GTFS Realtime data.
The Buffer component inputs extracted data $d_{i}$ and outputs sets $A$, $B$, $C$ taken from buffer indicies $d_{0},d_{h},$$d_{h+1},$ respectively. It uses RAM or “Memory” as a mechanism to store the contents and is controlled by $h$, as previously described in Computational Procedure. The purpose of this component is to roll the extracted data into and out of memory.
The Subset component inputs sets $A$, $B$, $C$ and outputs the realtime idling events $\mathcal{Y}$. It uses a series of set operations or “Operators” as a mechanism, and is controlled by $m$, as previously described in Computational Procedure. The purpose of this component is to compute the primary logic of Algorithm 1 GRD-TRT-BUF-4I.
Finally, the Store component inputs the realtime idling events $\mathcal{Y}$ and outputs them as historical information “Data”. It uses a database or “Disk” as a mechanism. The purpose of this component is to capture the historical record of realtime idling events so that they can be queried in downstream analyses. No explicit controls are required.
[…]
Figure 5: IDEF0 System Architecture of GRD-TRT-BUF-4I Algorithm
2.5.2 System Design
The system design implements a more detailed construction of the System Architecture. It casts the IDEF0 architecture into a common Extract, Transform, Load (ETL) microservice design pattern based on geographic region (US East , US West , etc.), exhibited in Figure 6. The system uses templated ETL pipelines per region, where each microservice corresponds to the ETL phase that it is nested within. Each microservice is then duplicated an arbitrary number of times, increasing redundancy to avoid single points of failure.
For each geographic region (Region 1, Region 2, ..., Region $N$), the Extract phase asynchronously extracts data from the associated GTFS Realtime servers (Server 1, Server 2, ..., Server $n$), previously outlined in System Architecture. The Transform phase then computes GRD-TRT-BUF-4I from the extracted data, as previously mentioned in Computational Procedure. Finally, the Load phase uses the “Write” microservices to insert the realtime idling event data into the Events table within the “Store” microservice.
Modular decoupling of components into microservices allows users to directly access any phase of the ETL pipeline throughout any and all geographic regions. The design not only allows this, but strongly encourages it. To use the realtime idling event data, users access the “Subset” microservice directly, rather than through “Write” or “Store”. This is achieved through the Subset API’s and is explained later in Data Availability.
Note that while casting the IDEF0 architecture into an ETL microservice design, the “Buffer” components are integrated within the “Subset” microservices along with parameters r, h, and m as a practical design choice. The “Write” microservices are introduced to appropriately decouple their associated function per geographic region, whereas the alternative would require a tightly coupled design with either “Subset” or “Store”. In addition, the Agency table is persisted within “Store” and is explained in the following section Records.
ExtractTransformLoadUserServer 1Server 2Server nnExtractSubsetBufferWriteServer 1Server 2Server nnExtractSubsetBufferWriteServer 1Server 2Server nnExtractExtractSubsetBufferWriteStoreAgencyEventsRegion 1Region 2Region NN
Figure 6: ETL Microservice System Design of IDEF0 Architecture
3 Records
There are three ways to use the recorded data. The first is through the Realtime Responses. The second is through a historical record of realtime responses in the Relational Database. The third is through the Static Files queried from the Relational Database.
3.1 Realtime Responses
The realtime responses are JavaScript Object Notation (JSON) objects generated by the Subset microservice. Connecting to any of the Subset API’s requires access to the region’s REST API endpoint via Websocket. This is explained later in Data Availability.
The JSON response will typically contain an array of key-value pairs. In the rare case that no idling events are detected, an empty response is returned without error. A typical JSON response contains eight keys: iata_id, vehicle_id, route_id, trip_id, latitude, longitude, datetime, and duration.
The JSON keys are equivalent to the fields found in Events Table. Each field is described in Table 15 in the following section Relational Database. An example of a typical JSON response is exhibited in Figure 7. Recall that this is $\mathcal{Y}$ from Subsetting Procedure.
[ { "iata_id": "NYC", "vehicle_id": "MTA NYCT_9750", "route_id": "M42", "trip_id": "MQ_D3-Weekday-SDon-012900_M42_301", "latitude": 40.7625617980957, "longitude": -74.00098419189453, "datetime": 1697178720, "duration": 90 }, { "iata_id": "NYC", "vehicle_id": "MTA NYCT_9890", "route_id": "M104", "trip_id": "MV_D3-Weekday-SDon-011000_M104_101", "latitude": 40.814937591552734, "longitude": -73.95511627197266, "datetime": 1697178722, "duration": 120 }, { "iata_id": "NYC", "vehicle_id": "MTA NYCT_5975", "route_id": "BX9", "trip_id": "KB_D3-Weekday-SDon-011000_BX9_602", "latitude": 40.84089279174805, "longitude": -73.87944030761719, "datetime": 1697178721, "duration": 60 } ]
Figure 7: Example of Realtime Websocket Response as JSON Object
3.2 Relational Database
The relational database is based on a common Dimensional Fact Model (DFM) implemented in PostgreSQL. It stores two tables. The first is the Agency table, which acts as the “dimension table”, containing static and descriptive attributes. The second is the Events table, which acts as the “fact table” that stores the realtime idling events. The two are joined on the common field iata_id or “natural key” as the primary key, exhibited in Figure 8.
Agency iata_id
agency
city country region continent
Events iata_id
vehicle_id
route_id trip_id latitude longitude datetime duration
Figure 8: DFM Relational Schema Between Agency Table and Events Table
3.2.1 Agency Table
The Agency table contains six fields: iata_id, agency, city, country, region, continent. Table 13 describes each field in greater detail. Table 14 provides a truncated example taken from the region, US West. Recall that this is the “dimension table” in the DFM, previously mentioned in Relational Database.
Field Description
iata_id IATA identifier. A unique three-letter code designating the fleet’s location.
agency The transit agency responsible for administering and operating the fleet.
city The name of the city associated with the IATA identifier and transit agency.
country The country where the city is located.
region The geographic region where the city is located.
continent The continent on which the country is located.
Table 13: Field Descriptions of Agency Table
iata_id agency city country region continent
LAX (Metro) Los Angeles United States United States West North America
SFO (MTC) San Francisco United States United States West North America
Table 14: Example of Agency Table from US West
3.2.2 Events Table
The Events table contains eight fields: iata_id, vehicle_id, route_id, trip_id, latitude, longitude, datetime, duration. They are the same as the keys found in the JSON object, previously mentioned in Realtime Responses. Table 15 describes each field in greater detail. Table 16 provides a truncated example taken from the geographic region, US West as an example. Recall that this is the “fact table” in the DFM, previously mentioned in Relational Database.
Field Description
iata_id IATA identifier. A unique three-letter code designates the fleet’s location.
vehicle_id Non-standard identifier of a single vehicle within a fleet.
route_id Non-standard identifier of each vehicle’s specified transit route.
trip_id Non-standard identifier of each vehicle’s on-network trajectory.
latitude Geocoordinate of the north–south position of an idling event (WGS84).
longitude Geocoordinate of the east-west position of an idling event (WGS84).
datetime Unix epoch timestamp designating the start date and time of an idling event.
duration Number of seconds elapsed since the start of an idling event.
Table 15: Field Descriptions of Events Table
iata_id vehicle_id route_id trip_id latitude longitude datetime duration
LAX 8538 70016… 16-13… 34.08… -118.38… 1704067086 80
SFO 8742 SF:11… SF:23… 37.73… -122.43… 1704067076 76
SEA 7235 53452… 10025… 47.72… -122.29… 1704066915 79
Table 16: Example of Events Table from US West
3.3 Static Files
There are three static files that can be downloaded in Comma Separated Value (.csv) format. Each file is a result from a table join between the Agency table and the Events table. Static files contain all fields found in both tables over a 24-hour collection period. The first file test-data-a.csv was collected from December 31, 2023 00:01:30 UTC to January 1, 2024 00:01:30 UTC. The second file test-data-b.csv was collected from January 4, 2024 01:30:30 UTC to January 5, 2024 01:30:30 UTC. The third file test-data-c.csv was collected from January 10, 2024 16:05:30 UTC to January 11, 2024 16:05:30 UTC. They can be downloaded at: https://doi.org/10.6084/m9.figshare.25224224 [81].
Figure 9 exhibits the query used to create the first static file from the relational database, previously mentioned in Relational Database. The subsequent files were created using this query corresponding to their respective datetime ranges. Table 17 provides a truncated example of the files. Again, they can be downloaded in .csv format. They are the same .csv files used for evaluating data quality in the following section Validation.
SELECT agency.*, events.vehicle_id, events.trip_id, events.route_id, events.latitude, events.longitude, events.datetime, events.duration FROM agency LEFT JOIN events ON agency.iata_id = events.iata_id WHERE events.datetime BETWEEN EXTRACT(EPOCH FROM TIMESTAMP ’2023-12-3100:01:30’) AND EXTRACT(EPOCH FROM TIMESTAMP ’2024-01-0100:01:30’);
Figure 9: PostgreSQL Query for Retrieving Static File(s)
iat.. age.. cit.. cou.. veh.. rou.. tri.. lat.. lon.. dat.. dur..
MIA Mia.. Mia.. Uni.. 21174 5639655 28437 25.6… -80.4… 1704066908 240
SYD Tra.. Syd.. Aus.. 43054… 2036432 2459… -33.8… 151.1… 1704066908 127
BNE Tra.. Bri.. Aus.. 73DD… 261484… 777-… -28.1… 153.5… 1704066908 90
Table 17: Example of Static File
4 Validation
Realtime data quality was assessed using three separate 24-hour validation periods. As previously mentioned in Static Files, the first was conducted from December 31, 2023 00:01:30 UTC to January 1, 2024 00:01:30 UTC using test-data-a.csv. The second from January 4, 2024 01:30:30 UTC to January 5, 2024 01:30:30 UTC using test-data-b.csv. The third from January 10, 2024 16:05:30 UTC to January 11, 2024 16:05:30 UTC using test-data-c.csv.
All three validation periods were arbitrarily specified and enforced a 72-hour minimum resting period to proxy independence. A battery of 113 individual tests were conducted for each 24-hour validation period to ensure realtime data quality throughout six overlapping categories: Data Types, Duplication, Missingness, Geolocation Error, Temporal Contiguity, and Expected Duration. Each category and its corresponding tests are explained in greater detail below.
4.1 Data Types
All data type test results were as expected for all validation periods. Exhibited in Table 18, test number 1 found the expected data object. Test numbers 2 to 7 found obvious string types - tests 8 to 10, also string types. Note that vehicle_id, route_id, and trip_id contained numerical values in some cases, but were found to be correctly encoded as string types. Geocoordinates from test numbers 12 and 13 were correctly encoded as float types. Test number 13 for datetime was correctly encoded as an integer type, as well as duration as an integer type in test number 14.
No. Field test-data-a.csv test-data-b.csv test-data-c.csv
1 Object DataFrame DataFrame DataFrame
2 iata_id string string string
3 agency string string string
4 city string string string
5 country string string string
6 region string string string
7 continent string string string
8 vehicle_id string string string
9 route_id string string string
10 trip_id string string string
11 latitude float float float
12 longitude float float float
13 datetime integer integer integer
14 duration integer integer integer
Table 18: Global Data Types by Field
4.2 Duplication
Of the 192,237 observations in test-data-a.csv, 197,411 observations in test-data-b.csv, and 209,329 observations in test-data-c.csv, none of them contained duplicate observations, nor did they contain duplicate fields. Although Data Types previously implied that all validation periods contained the same unique fields, Table 19 exhibits this explicitly from the results of test number 15. In addition to all unique fields, test number 16 found all unique observations.
No. Dimension test-data-a.csv test-data-b.csv test-data-c.csv
15 Fields 0.00 % 0.00 % 0.00 %
16 Observations 0.00 % 0.00 % 0.00 %
Table 19: Global Percentages of Duplication by Dimension
4.3 Missingness
Two of thirteen fields contained missing values across all validation periods. In Table 20, test number 24 found that route_id averaged (unweighted) 7.66% missingness. Test number 25 found that trip_id averaged (unweighted) 1.38% missingness. Although both route_id and trip_id contained missing values, when tested together, no missing values were present. Test number 26 found that every observation contained either a route_id or trip_id or both.
No. Field test-data-a.csv test-data-b.csv test-data-c.csv
17 iata_id 0.00 % 0.00 % 0.00 %
18 agency 0.00 % 0.00 % 0.00 %
19 city 0.00 % 0.00 % 0.00 %
20 country 0.00 % 0.00 % 0.00 %
21 region 0.00 % 0.00 % 0.00 %
22 continent 0.00 % 0.00 % 0.00 %
23 vehicle_id 0.00 % 0.00 % 0.00 %
24 route_id 7.50 % 8.67 % 6.82 %
25 trip_id 1.64 % 1.31 % 1.20 %
26 route_id | trip_id 0.00 % 0.00 % 0.00 %
27 latitude 0.00 % 0.00 % 0.00 %
28 longitude 0.00 % 0.00 % 0.00 %
29 datetime 0.00 % 0.00 % 0.00 %
30 duration 0.00 % 0.00 % 0.00 %
Table 20: Global Percentages of Missingness by Field(s)
4.4 Geolocation Error
4.4.1 General Error
General geolocation tests did not find any value errors across all validation periods. Table 21 and Table 22 exhibit the test results in the latitude and longitude fields, respectively. Tests 31 to 33 corresponding to latitude and tests 34 to 36 corresponding to longitude, did not find any zero values or those outside the projected bounds of the World Geodetic System 1984 (WGS84). Recall tests 11 and 12 in Data Types, and 27 and 28 in Missingness, also did not contain value errors.
No. latitude test-data-a.csv test-data-b.csv test-data-c.csv
31 Zero Values 0.00 % 0.00 % 0.00 %
32 Exceed Max. (> $\text{ }90^{\circ}$) 0.00 % 0.00 % 0.00 %
33 Exceed Min. ($90^{\circ}$) 0.00 % 0.00 % 0.00 %
Table 21: Global Percentages of Erroneous Values in latitude Field (WGS84)
No. longitude test-data-a.csv test-data-b.csv test-data-c.csv
34 Zero Values 0.00 % 0.00 % 0.00 %
35 Exceed Max. (> $\text{ }180^{\circ}$) 0.00 % 0.00 % 0.00 %
36 Exceed Min. ($180^{\circ}$) 0.00 % 0.00 % 0.00 %
Table 22: Global Percentages of Erroneous Values in longitude Field (WGS84)
4.4.2 Spatial Point Error
In addition to the value error tests previously mentioned in General Error, each idling event was tested to ensure the recorded geolocation was within a reasonable proximity to the shape of the corresponding route path. The shape of the route paths and their identifiers were collected from the GTFS static data corresponding to the GTFS Realtime source. The shape of the route paths were used as a geographic benchmark to validate the measured geolocation of a given idling event - anticipating the route_id or trip_id should be geographically near the shape of its corresponding route path. Recall test 26 from Missingness found that every observation contained either a route_id, trip_id, or both.
The notion of reasonable proximity was operationalized as a distance threshold specified with three primary considerations. First, the width of a typical urban arterial street was used as a tolerable margin of error between approximately 15 to 20 meters throughout varying degrees of urban density [82, 83, 84]. Second, Global Positioning System (GPS) location errors were estimated between 5 to 10 meters, especially in the context of urban environments where buildings, infrastructure, etc. can all obstruct reported vehicle locations [85, 86, 87]. Third, a preliminary test was drawn on a series of distance thresholds that combined the first two considerations and specified its midpoint to ensure that 25 meters was flexible enough to afford meaningful results, as exhibited in Figure 10.
0101020203030404050506060707080809090100100010102020303040405050606070708080909010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (a): test-data-a.csvUnweightedWeighted
0101020203030404050506060707080809090100100010102020303040405050606070708080909010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (b): test-data-b.csvUnweightedWeighted
0101020203030404050506060707080809090100100010102020303040405050606070708080909010010025Distance Threshold $D_{m}$ (meters)Average Spatial Point Error (%)Figure 10 (c): test-data-c.csvUnweightedWeighted
Figure 10: Global Average Spatial Point Error Relationship to Distance Threshold $D_{m}$
Spatial point error testing begins with the global distance threshold $D_{m}$ specified in meters:
$D_{m}=25\text{m}$ (20)
$D_{d}$ is then computed from $D_{m}$ by converting meters to degrees, using radians:
$D_{d}=\text{degrees}/\text{m}=1/111.32\times 1000\times\cos\left(\phi\times\frac{\pi}{180}\right)$ (21)
where $\phi$ is the average latitude of a given city in degrees:
$\phi=\frac{1}{n}\sum^{n}_{i=1}\text{latitude}_{i}$ (22)
A k-Dimensional Tree (K-D Tree) [88] is constructed for the shape of each route path $s$ in the set of all route path shapes $S$ of a given city, composed of a set of geocoordinates $\{p_{1},p_{2},\ldots,p_{m}\}$, where each geocoordinate $p_{i}$ contains a tuple $(\text{latitude}_{i},\text{longitude}_{i})$, such that:
 $\forall\text{ }s\in S:\text{K-D Tree}_{s}=KDTree(\{(\text{latitude}_{i},\text{longitude}_{i})\mid p_{i}\in s\})$ (23)
Given a set of idling events $P$, where each idling event $P_{j}\in P$ contains a tuple $(\text{latitude}_{j},\text{longitude}_{j})$, a given $P_{j}$ within $D_{d}$ of $s$ is computed as:
$C(P_{j},D_{d})=\sum_{s\in S}\textbf{1}\left(dist(P_{j},s)\leq D_{d}\right)$ (24)
where $C(P_{j},D_{d})$ is a counter of $P_{j}$ and $D_{d}$, 1 is an indicator function, and $dist(P_{j},s)$ is the distance between $P_{j}$ and $s$ using $\text{K-D Tree}_{s}$.
Finally the spatial point error $e$ is computed as the percentage of idling events $P_{j}$ that are outside the bounds of the global distance threshold $D_{d}$ from all route path shapes $S$ of a given city:
$e=1-\left(\frac{\sum_{j=1}^{n}C(P_{j},D_{d})}{n}\right)$ (25)
If $e$ is 0%, all idling events were within the distance threshold corresponding to the shape of the route path. Any $e$ greater than 0% indicates that there were idling events measured outside of it.
Table 23 exhibits the global average spatial point errors for all validation periods. The reported averages are both weighted and unweighted. Weighted averages are weighted by the number of observations per city, whereas the unweighted averages assign equal weight to all cities within the context of how they are reported, globally or regionally. Table 24 to Table 34 exhibit the test results for individual cities and their regional unweighted and weighted averages.
Globally, test numbers 37 and 38 found the unweighted and weighted average error did not exceed 6.50% and 5.05%, respectively. That is to say, the global average spatial point accuracy of measured idling events was at least 93.50%. Regionally, test number 52 in Table 25 found that 11.06% was the highest unweighted average error. Test number 67 in Table 27 found that 12.29% was the highest weighted average error. In other words, the average spatial point accuracy of measured idling events for any given region was at least 87.71%.
Note that in Table 31, Table 32, and Table 34, cities without results encountered preliminary test errors. In those cases, identity mappings could not be established between route_id or trip_id and the shape of the corresponding route path, either because of unmatched identifiers or the shape of the corresponding route path was unavailable. In successful preliminary tests, the city is indicated along with the GTFS static data source and the geolocation error.
No. Global test-data-a.csv test-data-b.csv test-data-c.csv
37 Average (Unweighted) 5.73 % 6.50 % 5.96 %
38 Average (Weighted) 5.05 % 4.67 % 5.00 %
Table 23: Global Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
39 New York GTFS [89] 2.93 % 2.97 % 2.87 %
40 Philadelphia GTFS [90] 6.72 % 8.16 % 6.39 %
41 Wash. D.C. GTFS [91] 5.24 % 5.60 % 5.97 %
42 Boston GTFS [92] 9.78 % 7.84 % 7.31 %
43 Pittsburgh GTFS [93] 2.49 % 7.95 % 6.91 %
44 Average (Unweighted) 5.43 % 6.50 % 5.89 %
45 Average (Weighted) 4.24 % 3.95 % 3.79 %
Table 24: US East Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
46 Los Angeles GTFS [94] 11.75 % 11.36 % 12.76 %
47 San Fran. GTFS [95] 5.29 % 5.13 % 5.26 %
48 San Diego GTFS [96] 3.91 % 5.99 % 6.79 %
49 Seattle GTFS [97] 2.10 % 2.49 % 1.91 %
50 Sacrameto GTFS [98] 6.71 % 35.04 % 5.11 %
51 Portland GTFS [99] 8.11 % 6.35 % 8.80 %
52 Average (Unweighted) 6.31 % 11.06 % 6.77 %
53 Average (Weighted) 6.17 % 5.72 % 6.27 %
Table 25: US West Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
54 Atlanta GTFS [100] 5.29 % 8.60 % 7.54 %
55 Miami GTFS [101] 8.89 % 9.90 % 12.30 %
56 Tampa GTFS [102] 3.91 % 0.00 % 0.00 %
57 Louisville GTFS [103] 9.83 % 18.42 % 15.96 %
58 Nashville GTFS [104] 10.34 % 10.35 % 13.87 %
59 Average (Unweighted) 6.87 % 9.45 % 9.93 %
60 Average (Weighted) 7.53 % 9.73 % 11.08 %
Table 26: US South Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
61 Minneapolis GTFS [105] 9.08 % 12.35 % 10.91 %
62 St. Louis GTFS [106] 0.60 % 4.22 % 1.42 %
63 Madison GTFS [107] 13.81 % 14.59 % 16.95 %
64 Columbus GTFS [108] 11.35 % 6.09 % 6.24 %
65 Des Moines GTFS [109] 4.98 % 4.13 % 2.49 %
66 Average (Unweighted) 7.96 % 8.28 % 7.60 %
67 Average (Weighted) 9.47 % 12.29 % 10.63 %
Table 27: US Central Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
68 Denver GTFS [110] 2.85 % 4.16 % 4.14 %
69 Phoenix GTFS [111] 10.26 % 11.26 % 12.03 %
70 San Antonio GTFS [112] 1.32 % 3.13 % 2.25 %
71 Austin GTFS [113] 0.74 % 0.11 % 0.25 %
72 Billings GTFS [114] —— % —— % —— %
73 Average (Unweighted) 3.79 % 4.66 % 4.67 %
74 Average (Weighted) 8.34 % 9.99 % 10.85 %
Table 28: US Mountain Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
75 Montréal GTFS [115] 1.71 % 1.93% 2.83%
76 York GTFS [58] 3.13 % 3.04 % 3.85 %
77 Hamilton GTFS [116] 17.54 % 14.95 % 16.41 %
78 Halifax GTFS [117] 1.09 % 1.08 % 0.89 %
79 Thunder Bay GTFS [118] 4.20 % 1.72 % 2.26 %
80 Average (Unweighted) 5.53 % 4.54 % 5.25 %
81 Average (Weighted) 2.45 % 2.40 % 3.31 %
Table 29: Canada East Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
82 Vancouver GTFS [119] 3.24 % 3.84 % 3.77 %
83 Calgary GTFS [120] 9.82 % 0.03 % 0.25 %
84 Edmonton GTFS [121] 0.98 % 1.34 % 4.41 %
85 Saskatoon GTFS [122] 4.38 % 5.58 % 9.47 %
86 Average (Unweighted) 4.60 % 2.70 % 4.47 %
87 Average (Weighted) 2.14 % 2.40 % 4.36 %
Table 30: Canada West Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
88 Amsterdam GTFS [123] 3.68 % 3.7% 3.39 %
89 Stockholm —— —— % —— % —— %
90 Helsinki GTFS [124] 2.46 % 2.95 % 1.79 %
91 Dublin GTFS [125] 10.61 % 2.90 % 3.61 %
92 Rome GTFS [70] 7.28 % 6.36 % 1.45 %
93 Average (Unweighted) 6.01 % 3.98 % 2.56 %
94 Average (Weighted) 4.76 % 3.44 % 3.05 %
Table 31: EU West Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
95 Warsaw —— —— % —— % —— %
96 Kraków —— —— % —— % —— %
97 Gdańsk —— —— % —— % —— %
98 Prague GTFS [126] 2.72 % 6.00 % 6.40 %
99 Average (Unweighted) 2.72 % 6.00 % 6.40 %
100 Average (Weighted) 2.72 % 6.00 % 6.40 %
Table 32: EU Central Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
95 Sydney GTFS [127] 1.48 % 1.63 % 1.83 %
96 Brisbane GTFS [128] 4.05 % 5.89 % 6.41 %
97 Adelaide GTFS [129] 9.77 % 7.21 % 6.96 %
98 Auckland GTFS [130] 7.67 % 7.70 % 7.67 %
99 Christchurch GTFS [131] 1.83 % 2.10 % 2.03 %
100 Average (Unweighted) 4.96 % 4.91 % 4.98 %
101 Average (Weighted) 5.79 % 6.01 % 6.45 %
Table 33: Oceania Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
No. City Routes test-data-a.csv test-data-b.csv test-data-c.csv
102 Dehli —— —— % —— % —— %
103 Average (Unweighted) —— % —— % —— %
104 Average (Weighted) —— % —— % —— %
Table 34: Asia Percentages of latitude – longitude 25m Outside GTFS Routes (WGS84)
4.5 Temporal Contiguity
Continuous validity and consistency was globally tested throughout all validation periods. Exhibited in Table 35, test numbers 105 and 106 did not find any invalid values in the datetime field. Test number 107 found the average (unweighted) proportion of time that no observations were measured for longer than 1 minute was 2.75%. Test number 108 found the average (unweighted) longest single interval that no observations were measured was 4 minutes, 24 seconds. Also, in test 109, the average (unweighted) elapsed time from the first observation to the last was 23 hours and 58 minutes within the context of a 24-hour maximum time horizon, exhibited in Figure 11.
No. datetime test-data-a.csv test-data-b.csv test-data-c.csv
105 Zero Values 0.00 % 0.00 % 0.00 %
106 Negative Values 0.00 % 0.00 % 0.00 %
107 Downtime (>1>1 min.) 2.16 % 2.56 % 3.53 %
108 Downtime Max. Interval 3 min. 57 sec. 5 min. 21 sec. 3 min. 56 sec.
109 Elapsed Time 24 hrs. 0 min. 23 hrs. 57 min. 23 hrs. 59 min.
Table 35: Global Temporal Contiguity in datetime Field
12345678910111213141516171819202122232402,0002{,}0004,0004{,}0006,0006{,}0008,0008{,}00010,00010{,}0001,0001{,}0003,0003{,}0005,0005{,}0007,0007{,}0009,0009{,} [^0] HourNumber of Observationstest-data-a.csvtest-data-b.csvtest-data-c.csv
Figure 11: Global Temporal Contiguity in datetime Field
4.6 Expected Duration
The global average idling duration was as expected throughout all validation periods. Exhibited in Table 36, test numbers 110 and 111 did not find any invalid values in the duration field. Test number 112 found when the minimum idling duration was adjusted to a commonly used policy definition longer than 5 minutes [8], the proportion of idling time denominated by total operational time was 38.74%. This is broadly consistent and within the 30% to 44% range from existing studies, previously mentioned in Introduction. Figure 12 exhibits this comparison, as well as when the minimum idling tolerance was unadjusted according to the parameterized definition $h=1$ minute.
When the minimum idling duration in test number 113 was unadjusted according to $h=1$ minute, the average (unweighted) proportion of idling increased to 56.96%. Although approximately 18% of the average proportion of idling increased as the minimum idling tolerance decreased from 5 minutes to 1 minute, it is concluded that the duration field is still reliable and valid, given its combined consistency with existing studies when adjusted to common policy definitions of idling. The material difference between tests 112 and 113 is likely not a measurement error, but rather a unique feature of this data that warrants further investigation in downstream analyses.
No. duration test-data-a.csv test-data-b.csv test-data-c.csv
110 Zero Values 0.00 % 0.00 % 0.00 %
111 Negative Values 0.00 % 0.00 % 0.00 %
112 Avg. Idle Adj. (>5>5 min.) 34.32 % 40.32 % 41.60 %
113 Avg. Idle (>h=1>h=1 min.) 52.48 % 58.62 % 59.78 %
Table 36: Global Expected Idling Duration Percentages in duration Field
0%30%44%100%Avg. Idle Adj.Avg. Idle48%62%53% 60%34% 42%Proportion of Idling
Figure 12: Global Expected Idling Duration Definitional Difference in duration Field
5 Data Availability
It is highly encouraged that this realtime data is indeed used in realtime applications and analyses, as well as with historical methods. In doing so, requires a functioning version of the software to be deployed. More information can be found below in Code Availability. In addition, updated versions of the Static Files are available at: https://doi.org/10.6084/m9.figshare.25224224 [81].
6 Code Availability
The repository containing the documentation, source code, configurations, and containers used as a part of this study is freely open and available on Github. The software is periodically maintained at the discretion of the authors and is licensed under General Public License Version 3 (GPLv3). It is available at: https://www.github.com/nickkunz/idling [132].
7 Conclusion
This is the first real-time detection system for urban transit bus idling measured on a global scale. Using GRD-TRT-BUF-4I, the system processes GTFS Realtime data from over 50 cities across North America, Europe, Oceania, and Asia, detecting approximately 200,000 idling events daily. A comprehensive data validation procedure was conducted using 113 individual tests across three 24-hour periods which demonstrated 93.50% spatial point accuracy and temporal contiguity exceeding 96%. The measured idling proportions were largely consistent with the existing literature when adjusted to common policy definitions of idling, confirming the reliability of the detection methodology.
Future work should focus on expanding geographic coverage, particularly in underrepresented cities and regions, as well as integrating additional vehicle telemetry data to enhance detection accuracy. The substantial difference between 1-minute and 5-minute idling thresholds observed in validation testing warrants further investigation to better understand short-duration idling patterns and their operational significance. As GTFS Realtime adoption continues to develop internationally, the system not only provides a practical implementation for detecting urban transit bus idling but also a conceptual framework for treating the issue as a coordinated global challenge rather than an isolated municipal concern.
8 Author Contributions
8.1 Authors and Affiliations
Systems Engineering, Cornell University
Nicholas Kunz & H. Oliver Gao
8.2 Contributions
Nicholas Kunz developed and implemented the methods, software, documentation, testing, and manuscript. H. Oliver Gao initiated and supervised the development of the investigation, manuscript, and resources for data storage and compute. Both authors collaborated equally to the concept. The authors have agreed to the published version of this manuscript.
9 Competing Interests
The authors declare no competing interests.
10 Acknowledgements
We would like to thank the transit agencies for providing public developer support for GTFS and GTFS Realtime and those that accommodated special requests for enhanced permissions and support during testing and development.
Thank you Dr. Mengjie Han at MTC (San Francisco), Andrew Lowe and Dante Avery at Swiftly (L.A., Miami, Tampa), NTA (Dublin), and Trafiklab (Stockholm). Also, thank you Dr. Graeme Troxell for your encouragement and preliminary feedback. This study was supported by Gao Labs and Cornell Systems Engineering.
11 Appendix
Algorithm 1 GRD-TRT-BUF-4I
Input: $r>0,h>0,m>0$
Output: $\mathcal{Y}=\bigl{\{}[v_{(i,j)\,T}]:v\in\mathcal{Y}\bigl{\}}$
1:$i=1$
2:while $i$ do
[…]
4:if $i\geq T=t+h+2$ then
5:  delete $d_{i+h+2}$ // 5: limit buffer length
6:  $A\leftarrow d_{0}$, $B\leftarrow d_{h}$, $C\leftarrow d_{h+1}$
7:  $n(A^{a}\cap B^{b})=A^{a}\cap B^{b}$
8:  $\mathcal{A}=\bigl{\{}[x_{(i,\,j=1,\ldots,5)\,t}]:x\in A^{n}\bigl{\}}$, $\mathcal{B}=\bigl{\{}[y_{(i,\,j=1,\ldots,5)\,t+h}]:y\in B^{n}\bigl{\}}$
9:  $\mathcal{H}\leftarrow\mathcal{H}_{i}=\mathcal{A}\cap\mathcal{B}$, $\mathcal{C}=\bigl{\{}[z_{(i,\,j=1,\ldots,5)\,t+h+1}]:z\in C\bigl{\}}$
10:  for $i\in\mathcal{H}$ do
11:   if $i\notin k_{i}$ then
12:    $k_{i}\leftarrow 0$
13:   end if
14:   if $i\notin\mathcal{C}$ then
15:    $k_{i}=k_{i}+1$
16:   else
17:    $k_{i}=0$
18:   end if
19:   if $k_{i}\geq m$ then
20:    delete $k_{i}$, $H_{i}$ // 20: bound append length
21:   end if
22:  end for
23:  $\mathcal{Y}=\mathcal{H}\cap\mathcal{C}$
24:  return $\mathcal{Y}$
25:  $\text{Store}\leftarrow\mathcal{Y}$
26:else
27:  continue
28:end if
29:$i=i+1$
30:pause $r$ sec // 30: limit request rate
31:end while